ABC, Inc. has issued a $1,000 par 6% annual coupon bond that is to mature in 20 years. If your required rate of return is 7.5%, what price would you be willing to pay for the bond?
Answer #1

Information provided:

Par value=future value= $1,000

Time= 20 years

Coupon rate= 6%

Coupon payment= 0.06*1,000= $60

Required return= 7.5%

The value of the bond today is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

N= 20

PMT= 60

I/Y= 7.5

Press the CPT key and PV to compute the present value.

The value obtained is 847.08.

Therefore, the value of the bond today is $847.08.

In case of any query, kindly comment on the solution.